Find (showing your reasoning) a formula for the term tn of the sequence defined by t1 = 4; tn = 2tn−1, n ≥ 2.


1
Expert's answer
2021-11-03T08:45:38-0400
"t_n=2t_{n-1},n\geq2"

Then

"r=\dfrac{t_n}{t_{n-1}}=\dfrac{2t_{n-1}}{t_{n-1}}=2, n\geq2"

This geometric progression with the initial term "t=t_1=4" and the common ratio "r=2"


"4, 4(2), 4(2)^2,..."

The nth term of this geometric progression is


"t_n=tr^{n-1}"

"t_n=4(2)^{n-1},n\geq1"